# Beyond the Request: Harnessing HTTP Response Headers for Cross-Browser Web Tracker Classification in an Imbalanced Setting

### v2

@@ 2.2. Other Detection Approaches
Focusing on JavaScript API calls, Wu et al. (Wu et al., 2016) developed DMTTrackerDetector, a system using decision trees in which observed API calls serve as features. Their system, comprising a 717-dimensional feature vector, achieved 97.8% accuracy and addressed obfuscation issues, showing the distinct JavaScript API usage by trackers and non-trackers.
@@ 5.1. Problem Formulation
  • (CH5) Data Similarity: The datasets collected across browsers can exhibit pronounced similarities in their feature distributions and values. Given the high dimensionality of our feature space, exhaustive pairwise comparisons can be computationally expensive and time-consuming, highlighting the challenge of accurately gauging data similarities in such a complex space.
@@ 5.2. Machine Learning Pipeline
Although many different resampling techniques exist to combat class imbalance, such as over- and undersampling (Hernandez et al., 2013), generative adversarial networks (Engelmann and Lessmann, 2021), or variational autoencoders (Fajardo et al., 2021), the question of whether this step is necessary remains. Some disadvantages of such methods are, for example, the removal of observations, the replication of existing observations, or the creation of synthetic samples. Selecting and applying a sampling method can lead to overfitting and over-optimism as well (Santos et al., 2018), notwithstanding the question of whether it is even feasible to train the models on a distribution that was not observed in real-world data. Addressing CH3, our pipeline preserves the observed distribution as the imbalance is mild  [^6] https://developers.google.com/machine-learning/data-prep/construct/sampling-splitting/imbalanced-data in the $\text{Chrome}_{22}$ dataset, i.e., a sufficient number of samples exists for each class, which allows proper training to differentiate between classes. We also compute metrics that account for or are unaffected by imbalance, and the selected ensemble models are inherently robust to this issue.
Lastly, following the findings of Niculescu-Mizil and Caruana (Niculescu-Mizil and Caruana, 2005), we apply Isotonic Regression instead of Platt Scaling to calibrate the probabilities calculated by our models. This is done by fitting an Isotonic Regression model to the models’ predictions as well as the true labels to adjust the predicted probabilities. The calibration function for each model $i$ is denoted as $C_{i}(f_{i}(x))$.
@@ 5.6. Evaluation Metrics
The first metrics that are commonly used in web privacy research are Accuracy, Precision, Recall, False Positives (FP), True Positives (TP), False Negatives (FN), True Negatives (TN), and F1-Score. Two accuracy scores can be calculated: one from the perspective of trackers, the positive label, and one from non-trackers, the negative label. However, accuracy is not useful with imbalanced datasets because a model could always predict the majority class and achieve accuracy with the same value as the distribution of said class. For this reason, we calculate the following additional metrics to quantify the performance of a classifier with an imbalanced dataset:
Balanced Accuracy:
Log-loss Score:
(3) $L_{log}(y,p)=-[y\cdot log(p)+(1-y)\cdot log(1-p)]$
where $y$ is the label class and $p$ is the probability.
Matthews-Correlation-Coefficient:
The Matthews-Correlation-Coefficient (MCC) (Matthews, 1975), a metric that has seen increasing popularity in recent years, is defined as:
AUPRC:
The last metric in our study is the Area Under the Precision-Recall Curve (AUPRC). This metric is relevant when dealing with imbalanced datasets. The Precision-Recall Curve is a plot that illustrates the trade-off between precision and recall for a classifier at various thresholds. The AUPRC is an aggregated measure of this curve, representing the classifier’s ability to correctly identify positive instances across different thresholds. In imbalanced settings, where one class is significantly underrepresented, AUPRC is a more informative and reliable metric than ROC-AUC, as it focuses on the performance of the classifier on the minority class, which is of interest for us (Saito and Rehmsmeier, 2015).
@@ 6. Experiments
In this section, we evaluate the effectiveness of our proposed methodology in detecting trackers. We begin by presenting our classifier results based on $\text{Chrome}_{22}$ data and follow up by testing them with data from $\text{Firefox}_{22}$ and $\text{Brave}_{22}$. Furthermore, we assess the longitudinal performance with the $\text{Chrome}_{23}$ dataset that was collected almost a year after $\text{Chrome}_{22}$. Note that this section only summarizes our results; performance metrics across all classifiers are reported in the Appendix 9.1 and 9.2.
@@ 6.2. Cross-browser and Longitudinal Performance
Our longitudinal performance analysis indicates a slight decline in the efficacy of the classifiers when transitioning from the $\text{Chrome}_{22}$ to the $\text{Chrome}_{23}$ dataset. We can observe a decrease in metrics such as AUC-ROC, Accuracy, and others. Notably, the MCC score between ET and XGBoost are getting similar. A trend which we can observe across most classifiers for $\text{Chrome}_{23}$ but also $\text{Firefox}_{22}$. However, the MCC scores of the gradient boosting models did not drop significantly more between $\text{Chrome}_{22}$ and $\text{Chrome}_{23}$, e.g., XGBoost went down from 0.838 to 0.823, whereas the ET suffered from 0.903 to 0.851. This suggests that while all models are affected by evolving data characteristics, ensemble methods, particularly some gradient boosting models, may offer a more stable predictive performance in longitudinal settings in contrast to the ET and RF. Nevertheless, ET and RF had better or similar performance compared to XGBoost. Applying isotonic calibration has only minor improvements compared to the uncalibrated models, as highlighted by Figure 9.
@@ 6.4. HTTP Request Header Performance
Figure 8. Comparison between our response-based, request-based, and T.EX-Graph’s Random Forest classifier. The response-based classifier outperforms its counterparts across all evaluated metrics.
@@ 7. Discussion
Spoofing of header values to circumvent our approach should not be a problem, as the values are not taken into account. However, this limits the models’ explainability in the sense that we lose the actual header values that might provide us with relevant information to understand trackers.
Our cross-browser approach reveals three things: firstly, the $\text{Chrome}_{22}$ dataset already captures most of the tracker structure. Secondly, due to the resulting effectiveness of a $\text{Chrome}_{22}$-trained classifier on out-of-distribution data, such as $\text{Firefox}_{22}$ and $\text{Chrome}_{23}$, it may not be necessary to re-train classifiers for these datasets. Thirdly, applying our approach to browsers like Brave that filter web tracking by default has a strong negative performance impact because the classifiers trained on unfiltered data are not able to effectively classify the modified structure of trackers in the filtered data.
@@
Utilizing the models for automated web tracker detection in real-life application scenarios is possible in two scenarios. First, HTTP response information is unavailable in an online (or live) detection scenario to block communication with web trackers. In this scenario, our model can enable or support reinforced learning of a complementary model on whose prediction blocking of HTTP communication is decided. This process, e.g., may imply manual inspection of HTTP communication where the predictions of our model are considered in cases where the primary model is wrong or indecisive. Second, as filter lists are still fundamental for web privacy protection, the manual process of generating filter rules can be optimized or outright replaced by utilizing our model. In this scenario, existing data sets or data from volunteers containing HTTP communication are used to identify web tracking requests. Filter rules for those can be dynamically generated and appended to conventional filter lists, which are already used.
@@
As one of the oldest methods for blocking trackers (Mazel et al., 2019), filter lists have proven to be reliable and effective. However, they have to be updated constantly, which is commonly done manually. A longitudinal study of nine years by Alrizah et al. (Alrizah et al., 2019) showed that human errors and delayed updates are a root cause for lower performance. Chen et al. (Chen et al., 2021) highlighted evasion methods to circumvent blockers and filter lists: (i) changing URLs, (ii) embedding code directly into the website with the script-tag, and (iii) combining functional and tracking code in one file. In summary, using filter lists for our ground truth is not without problems and may impact the evaluation of our classifiers. Nevertheless, they present a popular ground truth in research to label datasets for training ML models.
@@ 8. Conclusion
In this work, we train $\text{Chrome}_{22}$-based classifiers on a subset of response headers to evaluate their effectiveness and usefulness in web tracker detection while retaining the natural imbalance of trackers. Our two best classifiers RF and ET achieve scores ranging from 0.9 to 0.98. In addition, we measure the cross-browser effectiveness, showing that multiple $\text{Chrome}_{22}$-based classifiers achieve high AUPRC but moderate MCC values for $\text{Firefox}_{22}$. However, we also observe large performance losses for $\text{Brave}_{22}$.
The longitudinal performance of the selected ML models deteriorates only by small margins, which highlights their potential due to the changing capabilities of trackers (Fouad et al., 2018; Bujlow et al., 2017; Englehardt and Narayanan, 2016). Still, periodic re-training might be necessary if the performance decreases significantly.
We observe that response headers represent characteristics of web trackers which were previously under-explored by past research. They can be used to differentiate well between trackers and non-trackers and provide additional information on the nature of tracking-related entities. Our data-driven approach combined with binarization reveales new and confirms previously identified features. Applying the same methodology to request headers results in a weaker performance, thus strengthening the feasibility of using response headers for web tracker detection. Moreover, headers represent only a part of an HTTP message, and research has shown that the message bodies are also relevant when it comes to understanding and identifying trackers (Li et al., 2015; Guarino. et al., 2020). Emerging technologies like Large Language Models could be used to analyze bodies in an efficient and meaningful way.
In future research, we plan to expand on our approach of leveraging response headers for web tracker detection by considering header values in feature engineering and exploring the viability of deep-learning approaches based on response header data. Additionally, we intend to integrate our approach with existing solutions to develop a holistic web tracking detection system.

### v3

@@ 2.2. Other Detection Approaches
Focusing on JavaScript API calls, Wu et al. (Wu et al., 2016) developed DMTTrackerDetector, a system using decision trees in which observed API calls serve as features. Their system, comprising a 717-dimen- sional feature vector, achieved 97.8% accuracy and addressed obfuscation issues, showing the distinct JavaScript API usage by trackers and non-trackers.
@@ 4.1. Descriptive Analysis of Headers
[…]
[Frequency distribution of headers for each browser, indicating the frequency with which each header is present in the dataset.]In this frequency plot we can observe that only a few number of headers are prevalent and most headers are rarely seen in the wild.
@@
[…]
[Venn diagram of HTTP response headers.]Venn diagram of HTTP response headers in the examined datasets, revealing a core set of 5672 headers that are common across Chrome, Firefox, and Brave. Unique headers are present in each browser, although $\text{Chrome}_{23}$ has the highest number of unique headers, which may reflect browser-specific features or experimental headers.
@@
[…]
[ECDF plots for two selected response headers.]Row (A) presents the ECDFs for both Chrome datasets for the Content-Length header values. We set a cut-off value at 10,000 to highlight the core observation and more than half of the responses have a value below this threshold – $\tilde{\mu}_{T22}=62$, $\tilde{\mu}_{NT22}=8068$ and $\tilde{\mu}_{T23}=45$, $\tilde{\mu}_{NT23}=8021.5$. Row (B) shows how similar the values across $\text{Chrome}_{22}$, $\text{Firefox}_{22}$, and $\text{Brave}_{22}$ are for the X-XSS-Protection header.
@@ 5.1. Problem Formulation
  • (CH5) Data Similarity: The datasets collected across brow- sers can exhibit pronounced similarities in their feature distributions and values. Given the high dimensionality of our feature space, exhaustive pairwise comparisons can be computationally expensive and time-consuming, highlighting the challenge of accurately gauging data similarities in such a complex space.
@@
[…]
[The parameterized pipeline automatically processes the collected datasets and trains the classifiers.]The parameterized pipeline automatically processes the collected datasets and trains the classifiers. The pipeline starts with data collection and labeling. After the data is initially parsed, it is prepared and processed. Lastly, the models are trained and tested, and calibrated.
@@ 5.2. Machine Learning Pipeline
Although many different resampling techniques exist to combat class imbalance, such as over- and undersampling (Hernandez et al., 2013), generative adversarial networks (Engelmann and Lessmann, 2021), or variational autoencoders (Fajardo et al., 2021), the question of whether this step is necessary remains. Some disadvantages of such methods are, for example, the removal of observations, the replication of existing observations, or the creation of synthetic samples. Selecting and applying a sampling method can lead to overfitting and over-optimism as well (Santos et al., 2018), notwithstanding the question of whether it is even feasible to train the models on a distribution that was not observed in real-world data.
Addressing CH3, our pipeline preserves the observed distribution as the imbalance is mild  [^6] https://developers.google.com/machine-learning/data-prep/construct/sampling-splitting/imbalanced-data in the $\text{Chrome}_{22}$ dataset, i.e., a sufficient number of samples exists for each class, which allows proper training to differentiate between classes. We also compute metrics that account for or are unaffected by imbalance, and the selected ensemble models are inherently robust to this issue.
Last, following the findings of Niculescu-Mizil and Caruana (Niculescu-Mizil and Caruana, 2005), we apply Isotonic Regression instead of Platt Scaling to calibrate the probabilities calculated by our models. This is done by fitting an Isotonic Regression model to the models’ predictions as well as the true labels to adjust the predicted probabilities. The calibration function for each model $i$ is denoted as $C_{i}(f_{i}(x))$.
@@ 5.6. Evaluation Metrics
The first metrics that are commonly used in web privacy research are Accuracy, Precision, Recall, False Positives (FP), True Positives (TP), False Negatives (FN), True Negatives (TN), and F1-Score. Two accuracy scores can be calculated: one from the perspective of trackers, the positive label, and one from non-trackers, the negative label. However, accuracy is not useful with imbalanced datasets because a model could always predict the majority class and achieve accuracy with the same value as the distribution of said class. For this reason, we calculate the following additional metrics to quantify the performance of a classifier with an imbalanced dataset: (i) Balanced Accuracy (BACC), (ii) Log-loss Score, (iii) Matthews-Correlation-Coefficient (MCC), and (iv) Area Under the Precision-Recall Curve (AUPRC). A further explanation for each metric is reported in the Appendix B.
@@ 6. Experiments
[…]
[t-SNE plots across browsers to gauge data similarity.]t-SNE plots using a representative sample across all browsers to gauge data similarity. Separate clusters exist for trackers and non-trackers, but there is significant overlap between both classes. The results are separated by class to highlight differences between trackers and non-trackers.
In this section, we evaluate the effectiveness of our proposed metho- dology in detecting trackers. We begin by presenting our classifier results based on $\text{Chrome}_{22}$ data and follow up by testing them with data from $\text{Firefox}_{22}$ and $\text{Brave}_{22}$. Furthermore, we assess the longitudinal performance with the $\text{Chrome}_{23}$ dataset that was collected almost a year after $\text{Chrome}_{22}$. Note that this section only summarizes our results. Performance metrics across all classifiers are reported in the Appendix D.
@@ 6.1. Intra-browser Performance
[…]
[Radar charts of the LR, ET, and XGBoost classifier for cross-browser performance.]Radar charts plotting the performance of three selected classifiers across a subset of 13 metrics, illustrating that, while all models perform well on the in-distribution test set, the ET model exhibits more consistent performance across the various test sets, suggesting better generalizability. The models show a reduced performance on out-of-distribution test sets, indicating potential overfitting to the training data.
@@ 6.2. Cross-browser and Longitudinal Performance
Our longitudinal performance analysis indicates a slight decline in the efficacy of the classifiers when transitioning from the $\text{Chrome}_{22}$ to the $\text{Chrome}_{23}$ dataset. We can observe a decrease in metrics such as AUC-ROC, Accuracy, and others. Notably, the MCC score between ET and XGBoost are getting similar. A trend which we can observe across most classifiers for $\text{Chrome}_{23}$ but also $\text{Firefox}_{22}$. However, the MCC scores of the gradient boosting models did not drop significantly more between $\text{Chrome}_{22}$ and $\text{Chrome}_{23}$, e.g., XGBoost went down from 0.838 to 0.823, whereas the ET suffered from 0.903 to 0.851. This suggests that while all models are affected by evolving data characteristics, ensemble methods, particularly some gradient boosting models, may offer a more stable predictive performance in longitudinal settings in contrast to the ET and RF. Nevertheless, ET and RF had better or similar performance compared to XGBoost. Applying isotonic calibration has only minor improvements compared to the uncalibrated models, as highlighted by Figure 9. The exceptions are GNB and AdaBoost, where estimates have improved, e.g., the uncalibrated AdaBoost initially centered most predictions around 0.5.
@@ 6.3. Feature Importance
[…]
[Feature importance of Chrome-based classifiers using HTTP response headers.]The top ten features are presented in order of importance across different classifiers based on the in-distribution test set. The heatmap illustrates the normalized importance values for each feature and classifier. We calculated Permutation feature importance for LR and GNB but not for the others as we do not have any high cardinality features.
@@ 6.4. HTTP Request Header Performance
Figure 8. Comparison between our response-based, request-based, and T.EX-Graph’s RF classifier. The response-based classifier outperforms its counterparts across all evaluated metrics.
[…]
[The RF classifier outperforms the baselines.]Comparison between our response-based, request-based, and T.EX-Graph’s RF classifier. The response-based classifier outperforms its counterparts across all evaluated metrics.
@@ 7.1. Experimental Results
Our cross-browser approach reveals three things: firstly, the $\text{Chrome}_{22}$ dataset already captures most of the tracker structure. Secondly, due to the resulting effectiveness of a $\text{Chrome}_{22}$-trained classifier on out-of-distribution data, such as $\text{Firefox}_{22}$, it may not be necessary to re-train classifiers for these datasets. Thirdly, applying our approach to browsers like Brave that filter web tracking by default has a strong negative performance impact because the classifiers trained on unfiltered data are not able to effectively classify the modified structure of trackers in the filtered data.
@@
7.2. Potential Evasion Approaches
Spoofing of header values to evade our approach should not be a problem, as the values are not taken into account. However, this limits the models’ explainability in the sense that we lose the actual header values that might provide us with relevant information to understand trackers and whether some headers are required by trackers or not.
Siby et al. (Siby et al., 2022) argued that content features, such as URLs, are more vulnerable to obfuscation than structural features. This assertion extends to the content of headers as well. However, our classifiers are not concerned with the content of headers, but rather with their mere presence. Malicious actors could employ two different techniques in this case: (i) the addition of headers and (ii) the removal of headers – both could be done intentionally or randomly. The complete removal, however, could potentially disrupt the functionality of trackers or non-trackers. The presented classifiers could be further enhanced through the integration or the extension of structural and flow features. In the case of Siby et al., the inclusion of these features contributed to the robustness of AdGraph.
The authors (Siby et al., 2022) developed a threat model and considered various techniques that trackers might utilize to evade detection. Although it is conceivable that an individual tracker could obfuscate its identity by employing a particular set of response headers to mimic non-trackers, if all trackers were to adopt a uniform set of headers or employ a similar strategy, they would collectively form a distinct cluster as illustrated in Figure 5, thereby differentiating themselves from non-trackers.
Furthermore, high coordination between trackers would be necessary to employ said strategies but due to the heterogeneity of trackers and their owners, this level of coordination is unlikely. Nevertheless, a large impact could be achieved if the owner’s trackers are highly prevalent such as Alphabet’s.
Lastly, the similarity between trackers and non-trackers with respect to their feature vectors is already notably, emphasizing the classifier’s effectiveness in differentiating between these two groups. This suggests that the classifier is already highly proficient in identifying and separating trackers from non-trackers.
7.3. Deployment Scenarios
@@
Utilizing the models for automated web tracker detection in real-life applications is possible in two scenarios. First, HTTP response information is unavailable in an online (or live) detection scenario to block communication with web trackers. In this scenario, our model can enable or support reinforced learning of a complementary model on whose prediction blocking of HTTP communication is decided. This process, e.g., may imply manual inspection of HTTP communication where the predictions of our model are considered in cases where the primary model is wrong or indecisive. Second, as filter lists are still fundamental for web privacy protection, the manual process of generating filter rules can be optimized or outright replaced by utilizing our model. In this scenario, existing datasets or data from volunteers containing HTTP communication are used to identify web tracking requests. Filter rules for those can be dynamically generated and appended to conventional filter lists, which are already used.
7.4. Limitations of Using Filter Lists
@@
As one of the oldest methods for blocking trackers (Mazel et al., 2019), filter lists have proven to be reliable and effective. However, they have to be updated constantly, which is commonly done manually. A longitudinal study of nine years by Alrizah et al. (Alrizah et al., 2019) showed that human errors and delayed updates are a root cause for lower performance. Chen et al. (Chen et al., 2021) highlighted methods to evade blockers and filter lists: (i) changing URLs, (ii) embedding code directly into the website with the script-tag, and (iii) combining functional and tracking code in one file. In summary, using filter lists for our ground truth is not without problems and may impact the evaluation of our classifiers. Nevertheless, they present a popular ground truth in research to label datasets for training ML models.
@@ 8. Conclusion
In this work, we train $\text{Chrome}_{22}$-based classifiers on a subset of binarized response headers to evaluate their effectiveness and usefulness in web tracker detection while retaining the natural imbalance of trackers.
We observe that response headers represent characteristics of web trackers which were previously under-explored by past research. They can be used to differentiate well between trackers and non-trackers and provide additional information on the nature of tracking-related entities. Our data-driven approach combined with binarization reveals new and confirms previously identified features. Applying the same methodology to request headers results in a weaker performance, thus strengthening the feasibility of using response headers for web tracker detection.
Our two best classifiers RF and ET achieve scores ranging from 0.9 to 0.98. In addition, we measure the cross-browser effectiveness, showing that multiple $\text{Chrome}_{22}$-based classifiers achieve high AUPRC but moderate MCC values for $\text{Firefox}_{22}$. However, we also observe large performance losses for $\text{Brave}_{22}$. The longitudinal performance of the selected classifiers deteriorates only by small margins, which highlights their potential due to the changing capabilities of trackers (Fouad et al., 2018; Bujlow et al., 2017; Englehardt and Narayanan, 2016). Still, periodic re-training might be necessary if the performance decreases significantly.
In future research, we plan to expand on our approach of leveraging response headers for web tracker detection by considering header values in feature engineering and exploring the viability of deep-learning approaches based on response header data. Additionally, we intend to integrate our approach with existing solutions to develop a holistic web tracking detection system. Moreover, headers represent only a part of an HTTP message, and research has shown that the message bodies are also relevant when it comes to understanding and identifying trackers (Li et al., 2015; Guarino. et al., 2020). Emerging technologies like Large Language Models could be used to analyze bodies in an efficient and meaningful way.